OpenStudy (iiariana):

Solve negative 7 over 3, the whole multiplied by x minus 3 equals negative 52. 21 39 45 59

OpenStudy (anonymous):

What do you mean the whole thing? The only thing I can see that we would solve is x.

OpenStudy (lynfran):

\[-\frac{ 7 }{ 3 }(x-3)=-52\]

OpenStudy (lynfran):

?

OpenStudy (iiariana):

It's like this (-7/3) * -3 = -52

OpenStudy (iiariana):

Yes it's like what @LynFran did

OpenStudy (lynfran):

\[-\frac{ 7 }{ 3}x+7=-52\]\[-\frac{ 7 }{ 3 }x=-52-7\]

OpenStudy (lynfran):

now to find x,...divide both sides of the equation by \[-\frac{ 7 }{ 3 }\]

OpenStudy (iiariana):

Ok

OpenStudy (iiariana):

@LynFran I did it and got 39?